OpenStudy (anonymous):

Assume that the business makes a profit with probability 0.4 in the first year. For each year thereafter, the business makes a profit with probability 0.6 if it made a profit in the previous year, and with probability 0.2 if it did not make a profit in the previous year. What is the probability that the business makes a profit in exactly two of its first three years?

OpenStudy (anonymous):

@ayeshaafzal221 @dan815 @freckles

OpenStudy (anonymous):

Can i help?

OpenStudy (anonymous):

please!

OpenStudy (anonymous):

if you know how go for it @Redneckk__Lifee

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

Best Answer: 3 sequences of profit (P : PPx: 0.2*0.7*0.3 = .042 PxP: 0.2*0.3*0.4 = .024 xPP: 0.8*0.4*0.7 = .224 add to get ans of 0.29 <---------